### v1

# Not All Attention Heads Are What You Need: Refining CLIP’s Image Representation with Attention Ablation

Abstract
This paper studies the role of attention heads in CLIP’s image encoder. While CLIP has exhibited robust performance across diverse applications, we hypothesize that certain attention heads negatively affect final representations and that ablating them can improve performance in downstream tasks. To capitalize on this insight, we propose a simple yet effective method, called Attention Ablation Technique (AAT), to suppress the contribution of specific heads by manipulating attention weights. By integrating two alternative strategies tailored for different application scenarios, AAT systematically identifies and ablates detrimental attention heads to enhance representation quality. Experiments demonstrate that AAT consistently improves downstream task performance across various domains, boosting recall rate by up to 11.1% on CLIP-family models for cross-modal retrieval. The results highlight the potential of AAT to effectively refine large-scale vision-language models with virtually no increase in inference cost.
1 Introduction
As a pioneering large-scale vision-language model (VLM), CLIP [28] has garnered widespread attention for its simple yet effective design and impressive performance across a wide range of downstream tasks [16, 40, 25, 6]. While early research focused on improving CLIP through advancements in data [20], supervision [38], and architecture [16], recent studies have shifted toward analyzing its learned representations [11, 18, 1]. These efforts aim to uncover the intrinsic characteristics of CLIP’s representations, providing insights for further enhancement.
A recent study [11] investigates CLIP’s visual representations by decomposing them across attention heads, revealing a strong correlation between certain attention heads and specific visual concepts. By manually removing heads associated with spurious cues in the last four transformer layers, the CLIP’s visual representation is improved for a targeted zero-shot classification task [11]. This suggests that individual attention heads contribute uniquely to the output representation, act as property-specific “filters” for visual concepts. Building on this insight, we systematically extend the analysis to all attention heads in CLIP’s image encoder across diverse downstream tasks.
Trained on vast amounts of internet-sourced image-text pairs, the CLIP-family models may include attention heads that encode task-irrelevant signals—such as domain biases or spurious cues—that hinder downstream performance. Additionally, some heads might overfit to noise inherent in the uncurated training data. These detrimental heads can be distributed across different transformer layers in the image encoder. We hypothesize that ablating such heads could refine CLIP’s representations, thereby consistently improving performance across diverse downstream tasks.
To verify our hypothesis, we conduct preliminary experiments assessing the impact of individual attention heads. We find that ablating even a single detrimental head can enhance CLIP’s downstream performance, with further gains from ablating simple combinations of negatively impactful heads. Inspired by this, we develop an efficient Attention Ablation Technique (AAT) that refines CLIP representations by evicting groups of detrimental heads. AAT formulates the problem as a combinatorial optimization task, solved using either a Genetic Algorithm (GA) or Back-Propagation (BP).
The key difference between GA-based AAT (AAT-GA) and BP-based AAT (AAT-BP) lies in how they identify detrimental attention heads. AAT-GA uses a fitness function that measures the distance between positive image-text pairs and hard negatives, while AAT-BP introduces a lightweight training scheme with a learnable gating parameter for each head to determine its relevance. Both strategies yield remarkable improvements in multi-domain retrieval and classification tasks, with their distinct advantages making them suitable for different use cases.
AAT leaves CLIP’s model parameters unchanged, instead manipulating attention weights of selected heads, as shown in Figure 1. It suppresses image token weights to a low level while amplifying the class token weight as compensation. This ensures minimal image token contribution from ablated heads to the output representation. As an agile and versatile technique, AAT offers potential advantages in resource-constrained scenarios, such as limited compute or scarce high-quality data, while preserves strong generalization across diverse domains. Our main contributions are as follows:
  • We extensively explore the roles of attention heads in CLIP’s image encoder and show that ablating detrimental heads improves representation quality. To our knowledge, this is the first systematic study on refining CLIP’s outputs by probing attention head impacts.
  • We introduce AAT, a simple yet effective method for automatically identifying and ablating detrimental attention heads without altering model parameters. AAT offers two alternative strategies—GA and BP—tailored for different application scenarios.
  • AAT achieves consistent and significant improvements in diverse downstream tasks, delivering recall rate gains of up to 11.1% for retrieval, with little additional inference cost.
Figure 1: An illustration of AAT-improved CLIP for text-to-image retrieval. “A-head $i$” denotes the $i$-th head in MHA. With model parameters frozen, AAT suppresses the selected detrimental heads in the image encoder. Cross marks indicate ablated heads while check marks denote retained ones.
2 Related Work
The Interpretability of CLIP’s Representation
Recent studies explore CLIP’s representations from various perspectives. [26] analyzes entanglement between words and images. [31] finds a strong connection between the modality gap and loss local minima. [22] reports a semantic shift toward background regions. [39] improves image-text matching explainability via gradients and heatmaps. [7] finds that CLIP exhibits distinctive outlier features tied to ImageNet shift robustness. Closer to AAT, [3] boosts interpretability by converting dense embeddings into sparse, concept-based ones, and [11] reveals spatial localization and property-specific roles of attention heads.
Attention Weight Manipulation
The attention mechanism is central to transformer-based models [4], and manipulating it in pre-trained models has shown notable benefits. [37] calibrates weights toward hidden attention sinks in LLMs; [18] reformulates attention weights for segmentation; [17] reweights semantic tokens to refine CLIP’s text embeddings; and [24] amplifies image token weights to alleviate hallucinations in VLMs. AAT shares conceptual similarities with these heuristic methods but introduces a novel manipulation strategy grounded in systematic empirical exploration.
Attention Head Pruning
The idea behind AAT bears some resemblance to early studies on language models (LMs) [27, 33, 2], which identify redundancy among attention heads. These methods use pruning to expedite inference, typically at the cost of slight quality loss. While head pruning has been extensively explored in LMs, its exploration in VLMs remains limited.
Parameter-Efficient Finetuning
PEFT [14, 19, 13, 32] are widely used in resource-constrained scenarios by updating only a small subset of model parameters. The proposed AAT-BP shares some similarities with PEFT, as it introduces hundreds of learnable parameters. However, key differences remain: PEFT typically functions as a black box, with the role of each parameter unclear, whereas AAT-BP offers explicit interpretability of every parameter. Moreover, AAT-BP is more parameter-efficient, requiring orders of magnitude fewer parameters than typical PEFT approaches [9].
3 The Impact of Individual Attention Heads
In a preliminary experiment, we probe the impact of each individual attention head in CLIP’s image encoder. Employing the technique described in Section 4.1, we ablate one head at a time in a CLIP model and assess its performance on a retrieval task. For instance, in ViT-B-based CLIP with 144 heads across 12 layers, this grid search yields 144 results—some outperforming the original model, others not. The top 5 highest-performing configurations are ranked in Table 1, with full results in Appendix B. Notably, ablating even a single detrimental head can boost mean recall by up to 1.2%.
Building on this observation, we conduct an experiment that ablates a straightforward combination of all individually identified detrimental heads—those that improve recall when ablated alone. As shown in Table 2, this naive combination consistently boosts retrieval performance across multiple test subsets. Interestingly, although the heads are selected independently, their joint ablation yields further gains. This motivates the development of more effective strategies for optimal head selection.
indices 11-6 10-7 10-9 9-9 11-0 vanilla
mean-R 81.1 80.8 80.8 80.7 80.7 79.9
Table 1: Top 5 single-head ablation configurations for text-to-image retrieval on the COCO-CN $val$ set [21]. “m-n” demotes the $n$-th head in the $m$-th transformer layer. Evaluations are based on a ViT-B-based CLIP model [35], using mean-R (the average of R@1, R@5, and R@10) as the metric.
Models $val$ $test$ $all$
vanilla CLIP [35] 79.9 81.1 41.7
naive joint head ablation 82.5 82.6 43.2
Table 2: Comparison of mean-R for text-to-image retrieval on the COCO-CN $val$, $test$ and $all$ set.
4 Attention Ablation Technique
We introduce AAT in this section. Section 4.1 explains the method for ablating individual attention heads. Sections 4.2 and 4.3 present two alternative strategies—GA and BP—for identifying a globally optimal set of detrimental heads, with each strategy tailored to different application scenarios.
Figure 2: The left $5\times 5$ matrix represents the original attention weight matrix, while the right one shows the attention weights after manipulating. CLS and IMG denote the class token and an image token, respectively, with the input token sequence length being 5. Deeper blue indicates lower attention scores approaching zero, while deeper red represents higher scores closer to one.
4.1 Ablating the Attention Head
A transformer layer consists of a multi-head attention module (MHA) and a feed-forward network (FFN) [10]. An input token sequence of length $N$ is structured with a leading class token followed by $N-1$ image tokens. Let $x^{l}_{h}$ represent the input to the $h$-th attention head of the MHA in the $l$-th transformer layer. The output of each head in the MHA ${y}^{l}_{h}$ can be formulated as follows:
$\mathcal{A}^{l}_{h}$ $=Softmax(\frac{f_{q}(LN(x^{l}_{h}))\cdot f_{k}(LN(x^{l}_{h})^{T})}{\sqrt{d_{k}}}),$ (1)
${y}^{l}_{h}$ $=\mathcal{A}^{l}_{h}\cdot f_{v}(LN(x^{l}_{h}))$ (2)
where $f_{q}$, $f_{k}$, and $f_{v}$ denote the projection layers for queries, keys, and values, respectively. $LN$ refers to LayerNorm, $d_{k}$ is the embedding dimension of each head, and $\mathcal{A}^{l}_{h}$ denotes the attention weight matrix with dimensions $N\times N$.
As shown in the formulations above, the attention weight matrix $\mathcal{A}$ is generated by a softmax function, ensuring that each row sums to 1. The output of each attention head, ${y}$, is computed as the matrix product of $\mathcal{A}$ and the projected values. To ablate an head, we manipulate $\mathcal{A}$ to suppress the contribution of image tokens in the output representation. An illustrative example is provided in Figure 2. Specifically, this manipulation involves the following two steps:
  1. 1.
    Reduce the importance of image tokens: Adjust $\mathcal{A}$ as $\mathcal{A}[:,j]$leftarrow $\mathcal{A}[:,j]\times\beta$, where $j\in\{1,2,...,N-1\}$.
  2. 2.
    Reweight each row of $\mathcal{A}$: Normalize $\mathcal{A}$ by setting $\mathcal{A}[i,j]$leftarrow $\mathcal{A}[i,j]~{}/~{}\mathcal{A}_{i}$, where $\mathcal{A}_{i}=\sum_{j}\mathcal{A}[i,j]$ and $i,j\in\{0,1,...,N-1\}$. This ensures that each row still sums to 1.
We use a hyper-parameter $\beta$ to control the degree of suppression applied to the weights of image tokens. Unless otherwise noted, we set $\beta=0.1$. After applying the two manipulation steps, the manipulated attention weight matrix replaces the original, while all other operations in MHA remain unchanged. This enables refinement of the output representation without altering model parameters.
4.2 AAT with Genetic Algorithm
As discussed in Section 3, ablating detrimental attention heads in CLIP can refine its output representation. Our objective is to identify an optimal selection of such heads. We formulate this as a combinatorial optimization problem and address it using a small validation set $\mathcal{D}$ containing a limited number of image-text pairs. We use a Genetic Algorithm (GA) [12], which evolves candidate solutions over generations to navigate the complex search space and acquire high-quality solutions.
Specifically, we represent all attention heads in the image encoder using a binary vector, where each element corresponds to a head (1 for ablated, 0 for retained). Using the validation set $\mathcal{D}$, GA is applied to optimize this vector configuration. To align with CLIP’s training objective that brings matching image-text pairs closer while pushes apart mismatched ones, we design a fitness function $\mathcal{F}$ that reflects this property. $\mathcal{F}$ serves as the optimization target for GA and is defined as follows:
$\mathcal{F}$ $=\frac{1}{N}\sum_{i=1}^{N}(S_{pos}^{i}-max_{j\in\mathcal{H}^{i}}(S_{neg}^{i,j}))$ (3)
where $N$ is the number of image-text pairs in $\mathcal{D}$. $S_{pos}^{i}$ is cosine similarity between the $i$-th text and its ground truth image. $S_{neg}^{i,j}$ is cosine similarity between the $i$-th text and the $j$-th image from $\mathcal{H}^{i}$, an updatable set of hard negatives for the $i$-th text. We detail the construction of $\mathcal{H}^{i}$ below.
Initially, we evaluate the vanilla CLIP on $\mathcal{D}$ for text-to-image retrieval. For each text query $i$, the top $k_{1}$ non-matching images (i.e., false positives) are collected as hard negatives to form the basis of $\mathcal{H}^{i}$. To avoid overfitting to these cases, we augment $\mathcal{H}^{i}$ with $k_{2}$ randomly sampled non-matching images from $\mathcal{D}$. At the end of each GA generation, these $k_{2}$ random samples are refreshed with newly sampled ones. This dynamic $\mathcal{H}$ is crucial to AAT-GA, as it encourages larger distance margins between true matches and evolving hard negatives, while also considering newly emerging challenging cases.
In Eq. 3, the first term of $\mathcal{F}$ promotes similarity between positive image-text pairs, while the second penalizes similarity with the hardest negatives in $\mathcal{H}$. Together with standard crossover and mutation operations, $\mathcal{F}$ effectively guides GA toward an optimal or near-optimal head selection for ablation.
Application Scenarios
AAT-GA refines CLIP’s representations entirely at inference time, making it well-suited for resource-constrained scenarios. It can be deployed on low-power, inference-targeted edge devices with limited-precision compute capabilities, such as INT8/INT16 chips (e.g., Google Coral Edge TPU, Ethos-N78, Mythic M1108 AMP). Moreover, AAT-GA is effective in data-scarce scenarios, requiring only a small number of samples for optimization (see Section 5.3.2), whereas supervised finetuning (SFT) tends to overfit under such conditions (see Appendix F).
4.3 AAT with Back-Propagation
Since GA may require more inference trials with a poorly initialized population, and not all resource-constrained scenarios lack floating-point (FP) computation power, we propose an alternative strategy for identifying detrimental heads using gradients: Back-Propagation (BP) optimization. Unlike the uniform suppression factor $\beta$ used in AAT-GA, BP introduces an individual learnable parameter $\alpha_{i}$ for each probed head. We replace $\beta$ with a head-specific factor $\beta_{i}\in[0,1]$, which determines the degree to which the $i$-th head is ablated. $\beta_{i}$ is defined as follows:
$\beta_{i}$ $=\text{Sigmoid}(\tau\cdot\alpha_{i})$ (4)
where the temperature $\tau$ is set above 1 to sharpen the sigmoid output, encouraging $\beta_{i}$ to approach extreme values (0 or 1) during training. Empirically, we set $\tau=5.0$ and initialize each $\alpha_{i}$ to 1.0, yielding an initial $\beta_{i}$ close to 1. As in AAT-GA, we use the same validation set $\mathcal{D}$ to optimize the parameters ${\alpha_{i}}$, guided by the contrastive supervision objective used in CLIP.
Application Scenarios
While BP might get stuck in a local optimum, its fast convergence makes it a practical and efficient solution for real-world applications. Its low compute and data requirements enhance its applicability for resource-limited settings. In spirit, AAT-BP resembles a PEFT method, yet it includes several orders of magnitude fewer learnable parameters than typical PEFTs [14, 13]. We believe this extreme efficiency stems from a deep understanding of CLIP’s attention heads.
tasks text-to-image retrieval image-to-text retrieval
splits the test set the all set the test set the all set
size method R@1 R@5 R@10 mR R@1 R@5 R@10 mR R@1 R@5 R@10 mR R@1 R@5 R@10 mR
B vanilla 38.8 64.2 73.8 58.9 13.0 26.5 33.8 24.4 57.8 81.7 88.6 76.0 24.2 43.1 52.1 39.8
AAT-GA 41.3 66.7 76.6 61.5 14.0 28.3 35.9 26.1 60.3 82.7 89.3 77.4 26.0 45.4 54.4 41.9
AAT-BP 40.5 66.2 75.9 60.9 13.9 28.1 35.6 25.9 58.9 82.5 89.5 77.0 24.8 44.2 53.1 40.7
L vanilla 43.9 68.8 78.3 63.7 16.1 31.3 39.1 28.8 59.5 82.3 89.2 77.0 25.5 44.9 53.9 41.4
AAT-GA 46.3 71.3 80.8 66.1 17.4 33.3 41.4 30.7 62.8 84.2 90.2 79.1 27.6 47.9 56.8 44.1
AAT-BP 46.0 71.8 80.9 66.2 17.2 33.2 41.3 30.6 60.6 83.3 89.7 77.8 25.6 45.6 54.7 42.0
H vanilla 47.4 72.1 80.9 66.8 18.5 34.8 42.9 32.1 60.6 84.6 90.9 78.7 26.0 46.1 55.4 42.5
AAT-GA 49.3 74.1 82.7 68.7 19.5 36.5 44.7 33.6 63.8 85.8 91.5 80.4 28.2 48.7 58.1 45.0
AAT-BP 48.6 73.5 82.2 68.1 19.1 35.9 44.1 33.0 63.7 85.9 91.5 80.4 28.1 48.8 58.0 45.0
Table 3: Comparison on MS COCO for retrieval. R@k denotes recall rate at rank k, and mR (namely mean-R) denotes the average of R@1, R@5, and R@10. Model sizes are denoted as B (ViT-B), L (ViT-L), and H (ViT-H). Best results are highlighted in bold. No further elaboration in the following.
5 Experiments
5.1 Experimental Settings & Implementation Details
Models
We evaluate AAT on CLIP [28] and Chinese-CLIP [35], across ViT-B, ViT-L, and ViT-H. Specifically, we use OpenCLIP released models [15] trained on LAION-2B (a subset of LAION-5B [30]) and Chinese-CLIP released models trained on about 200M Chinese image-text pairs.
AAT Optimization Data
As introduced in Sections 4.2 and 4.3, we use a small validation set $\mathcal{D}$ to optimize AAT. For CLIP, we randomly sample 1k image-text pairs from the MS COCO train set as $\mathcal{D}$. For Chinese-CLIP, we directly use the COCO-CN val set that contains 1k image-text pairs.
AAT-GA Implementation
The key hyper-parameters for GA are as follows: the population size is set based on the number of attention heads—48 for ViT-B, 96 for ViT-L, and 128 for ViT-H. Two-point crossover is applied with a probability of 0.9, and mutation occurs with a probability of 0.5 using a flip-bit strategy. Tournament selection with a size of 3 is used. The evolution runs for up to 100 generations, with early stopping triggered by stagnant fitness gains or low population diversity.
AAT-BP Implementation
The settings for training $\alpha$ in BP are as follows: a learning rate of 2e-2 for CLIP and 5e-2 for Chinese-CLIP, with a batch size of 256. ViT-B-based models are trained for 32 epochs, and larger models for 64. Other hyper-parameters follow the OpenCLIP codebase [15].
Benchmarks
We benchmark AAT-improved CLIP on MS COCO [23], Flickr30k [36], and ReCoS [5], and evaluate AAT-improved Chinese-CLIP on two widely used Chinese retrieval datasets: COCO-CN [21] and Flickr30k-CNA [34]. For MS COCO, Flickr30k, COCO-CN, and Flickr30k-CNA, we create two evaluation splits: the test set (i.e., the original test split) and the all set (including all samples in the dataset). This setup enables a comprehensive analysis of AAT in large-scale retrieval. To further broaden the evaluation, we also assess AAT on the val set of ImageNet-1k [8] for zero-shot classification. Details on datasets and evaluation metrics are provided in Appendix A.
tasks text-to-image retrieval image-to-text retrieval
splits the test set the all set the test set the all set
size method R@1 R@5 R@10 mR R@1 R@5 R@10 mR R@1 R@5 R@10 mR R@1 R@5 R@10 mR
B vanilla 67.20 88.68 93.06 82.98 30.54 52.17 61.27 47.99 85.90 97.90 99.10 94.30 48.16 72.76 81.18 67.36
AAT-GA 70.26 90.06 94.42 84.91 32.64 54.90 64.04 50.53 87.20 97.10 98.90 94.40 50.47 74.09 81.89 68.82
AAT-BP 69.76 90.08 93.96 84.60 32.69 54.79 63.98 50.49 86.30 97.70 99.00 94.33 49.58 73.77 81.97 68.44
L vanilla 73.52 91.84 95.36 86.91 36.26 58.98 67.91 54.39 87.00 98.20 99.50 94.90 48.90 74.39 83.09 68.79
AAT-GA 75.74 92.72 95.92 88.13 38.28 61.16 69.93 56.46 88.10 98.50 99.40 95.33 52.68 76.96 84.90 71.51
AAT-BP 76.20 93.12 95.96 88.43 38.47 61.75 70.62 56.95 87.50 98.10 99.40 95.00 50.89 75.71 83.70 70.10
H vanilla 76.10 93.44 96.42 88.65 41.18 64.03 72.61 59.27 88.60 98.70 99.60 95.63 51.10 77.44 85.64 71.39
AAT-GA 77.84 94.50 96.94 89.76 43.10 66.20 74.51 61.27 90.40 98.90 99.60 96.30 54.98 80.06 87.51 74.18
AAT-BP 77.78 94.48 96.92 89.73 42.62 65.78 74.32 60.91 90.70 98.90 99.60 96.40 56.16 80.79 88.05 75.00
Table 4: Comparison on Flickr30k for retrieval.
types natural images non-natural images
tasks ACT CEL COL CNT FIG OBJ OCR POS SCE mean CR NC TT mean
vanilla 73.6 74.0 73.4 59.8 82.8 79.6 77.0 65.8 87.0 74.8 27.7 21.0 2.7 17.1
AAT-GA 75.8 77.7 76.8 60.4 86.4 81.4 81.3 67.4 89.0 77.4 20.3 21.3 2.3 14.6
AAT-BP 74.8 77.0 76.2 60.6 86.6 83.0 79.7 71.6 88.0 77.5 22.3 18.3 2.3 14.3
Table 5: Text-to-image retrieval on ReCoS based on ViT-B, evaluated by R@1.
5.2 Main Results
5.2.1 AAT-improved CLIP for Retrieval
We evaluate CLIP models improved using both AAT-GA and AAT-BP, reporting retrieval performance across different sizes, including ViT-B, ViT-L, and ViT-H. Results are presented for MS COCO, Flickr30k, and ReCoS, with comparisons against the baseline vanilla CLIP.
MS COCO
Table 3 presents the results on MS COCO, achieved by AAT-GA and AAT-BP for text-to-image and image-to-text retrieval. For text-to-image retrieval, GA and BP perform comparably, with mean-R differences under 1%. On the test set, AAT yields a 1.3$\sim$2.6gains in mean-R across model sizes, and over 1.5% on average for the all set. For image-to-text retrieval, AAT achieves a 0.8$\sim$2.1mean-R gain on the test set and 0.6$\sim$2.7on the all set.
Flickr30k
Table 7 reports results on Flickr30k, showing a 1.08$\sim$1.93mean-R improvement on the test set and 1.64$\sim$2.56on the all set for text-to-image retrieval using AAT-improved CLIP. Notably, gains on the test set tend to diminish at higher recall levels, likely due to saturation. A similar trend is observed for image-to-text retrieval as well, yet with consistent mean-R improvements of 0$\sim$0.8on the test set and 1.08$\sim$3.61on the all set.
ReCoS
To demonstrate the broader effectiveness of AAT, Table 5 provides text-to-image retrieval results on the challenging ReCoS, spanning 12 diverse subsets across various domains (see Appendix A for subset abbreviations). These subsets are grouped into two categories based on image domains: natural and non-natural. The natural includes real-world imagery, such as landscapes, animals, movie scenes, and profile photos. In contrast, the non-natural comprises synthetic visuals like code snippets, characters, and formulas, typically displayed on white or gray backgrounds.
AAT consistently outperforms on natural-scene subsets, with a mean R@1 gain of around 3%. However, its performance declines on non-natural images. We attribute this drop to two main factors: (1) significant domain shifts between the test images and those used to identify detrimental attention heads (see Appendix G); and (2) the inherent limitations of CLIP models. Tasks such as CR (code reasoning), NC (numerical calculation), and TT (text translation) demand beyond visual perception, but understanding and reasoning capabilities that CLIP is not explicitly trained for.
tasks text-to-image retrieval image-to-text retrieval
splits the test set the all set the test set the all set
size method R@1 R@5 R@10 mR R@1 R@5 R@10 mR R@1 R@5 R@10 mR R@1 R@5 R@10 mR
B vanilla 62.2 86.9 94.3 81.1 24.0 45.4 55.8 41.7 56.2 83.8 93.4 77.8 22.6 43.0 52.7 39.4
AAT-GA 64.9 89.9 96.5 83.8 27.6 50.0 60.4 46.0 62.6 89.0 95.3 82.3 25.5 47.1 57.2 43.3
AAT-BP 66.0 90.7 96.3 84.3 27.4 50.4 60.6 46.1 63.1 90.4 95.6 83.0 26.2 48.3 58.1 44.2
L vanilla 63.9 88.7 94.5 82.4 26.7 48.6 58.5 44.6 60.4 84.2 93.3 79.3 24.4 44.6 54.5 41.2
AAT-GA 66.5 91.4 96.1 84.7 30.1 52.9 63.2 48.7 65.9 89.0 95.4 83.4 29.8 51.6 61.6 47.7
AAT-BP 68.2 91.1 96.6 85.3 30.2 52.9 63.3 48.8 67.3 90.6 96.9 84.9 30.4 52.8 63.5 48.9
H vanilla 69.6 89.9 95.8 85.1 30.4 52.4 62.2 48.3 63.1 86.7 93.0 80.9 26.3 46.5 56.0 42.9
AAT-GA 72.3 92.3 96.5 87.0 33.8 56.8 66.6 52.4 66.6 89.0 94.5 83.4 30.0 51.1 60.4 47.1
AAT-BP 73.7 92.7 97.6 88.0 35.5 59.4 69.1 54.7 72.4 93.1 97.3 87.6 35.4 58.4 68.3 54.0
Table 6: Comparison on COCO-CN for retrieval.
tasks text-to-image retrieval image-to-text retrieval
splits the test set the all set the test set the all set
size method R@1 R@5 R@10 mR R@1 R@5 R@10 mR R@1 R@5 R@10 mR R@1 R@5 R@10 mR
B vanilla 62.12 86.62 92.52 80.42 24.35 44.78 54.36 41.16 73.70 93.20 97.00 87.97 33.45 56.84 66.56 52.28
AAT-GA 65.60 89.04 94.18 82.94 26.79 48.20 57.91 44.30 77.70 95.90 97.50 90.37 36.76 60.51 70.10 55.79
AAT-BP 66.58 89.28 94.36 83.41 27.68 49.51 59.17 45.45 78.60 95.50 97.90 90.67 37.18 61.46 71.07 56.57
L vanilla 67.58 89.52 94.28 83.79 28.79 50.39 59.80 46.33 80.00 96.50 98.20 91.57 37.15 61.59 71.05 56.59
AAT-GA 71.48 91.60 95.52 86.20 32.45 55.02 64.51 50.66 87.00 97.60 98.90 94.50 45.17 69.78 78.53 64.50
AAT-BP 70.90 91.34 95.70 86.00 31.69 54.46 64.10 50.08 87.50 97.40 99.10 94.67 44.25 69.05 77.82 63.71
H vanilla 70.94 91.30 95.30 85.85 33.92 56.27 65.38 51.86 81.40 97.00 98.90 92.43 42.35 66.47 75.43 61.42
AAT-GA 74.86 93.00 96.50 88.12 37.54 60.46 69.47 55.82 85.40 98.10 99.00 94.17 48.23 72.16 80.29 66.89
AAT-BP 75.30 93.40 96.78 88.49 37.64 60.68 69.64 55.99 88.60 97.50 99.30 95.13 52.12 75.71 83.38 70.40
Table 7: Comparison on Flickr30k-CNA for retrieval.
datasets ImageNet-1k MS COCO Flickr30k
sizes ViT-B ViT-L ViT-H ViT-B ViT-L ViT-H ViT-B ViT-L ViT-H
vanilla 68.12 72.62 76.17 58.9 63.7 66.8 82.98 86.91 88.65
AAT-GA 69.17 74.11 77.28 61.2 65.8 68.6 84.67 88.02 89.67
AAT-BP 69.01 73.76 77.25 61.0 65.6 68.2 84.42 88.34 89.61
Table 8: Top-1 accuracy on the ImageNet-1k val set for zero-shot classification, and mean-R on the test sets of MS COCO and Flickr30k for text-to-image retrieval.
5.2.2 AAT-improved Chinese-CLIP for Retrieval
To further validate the versatility of AAT, we evaluate it on non-English linguistic CLIP models. Following the same evaluation protocol as with CLIP, we assess AAT-improved Chinese-CLIP on COCO-CN and Flickr30k-CNA, comparing against their vanilla counterparts.
COCO-CN
The text-to-image retrieval results are shown in Table 6. On the test set, AAT achieves a 1.9$\sim$3.2higher mean-R across model sizes using both GA and BP. On the all set, the gain is even more pronounced, reaching up to 6.4%. For image-to-text retrieval, AAT delivers a 2.5$\sim$6.7improvement on the test set and 3.9$\sim$11.1on the all set. These advances surpass those observed in CLIP, suggesting its effectiveness is correlated with the quality of the original representations.
Flickr30k-CNA
As shown in Table 7, AAT improves mean-R by 2.21$\sim$3.2on the test set and 3.14$\sim$4.33on the all set for text-to-image retrieval. For image-to-text retrieval, it outperforms the baseline by around 3% on the test set and up to 8.98% on the all set. Consistent with the results on COCO-CN, AAT delivers greater improvements for Chinese-CLIP compared to CLIP.
5.2.3 AAT-improved CLIP for Zero-shot Classification
Since AAT is primarily designed for retrieval, we extend its application to zero-shot classification on ImageNet-1k, with minimal but essential adaptation. We identify a mismatch between the text templates used in ImageNet-1k classification (e.g., “a photo of a CLASS_NAME”) and the natural language captions in the original optimization set $\mathcal{D}$ (e.g., “An old woman sits on a bench”). To resolve this, we construct a new $\mathcal{D}$ by incorporating 500 image-text pairs from the ImageNet-1k training set that follow the zero-shot template style, and re-run AAT on CLIP using this adapted $\mathcal{D}$.
We evaluate the resulting models on the ImageNet-1k val set. As shown in Table 8, they consistently achieve 1$\sim$1.5gains in top-1 accuracy. To further verify AAT’s generalization, we test retrieval performance of these models on MS COCO and Flickr30k. The models continue to deliver strong improvements, with no noticeable degradation compared to counterparts optimized with the original $\mathcal{D}$ (see Table 3 and 4), with differences of less than 0.6% on MS COCO and 0.3% on Flickr30k.
Figure 3: Mean-R on the COCO-CN all set vs. $\beta$ values in AAT-GA, using the ViT-B-based model.
the size of $\mathcal{D}$   0   100   200   500  1000
   CLIP AAT-GA   24.4   25.4   25.7   26.0   26.1
AAT-BP   24.4   25.2   25.4   25.7   25.9
   Chinese-CLIP AAT-GA   41.7   45.1   45.3   45.9   46.0
AAT-BP   41.7   44.7   45.2   45.7   46.1
Table 9: Mean-R vs. the size of $\mathcal{D}$: AAT-improved CLIP is evaluated on the MS COCO all set, and AAT-improved Chinese-CLIP on the COCO-CN all set, both using ViT-B-based models.
splits the test set the all set
metric R@1 R@5 R@10 mean-R R@1 R@5 R@10 mean-R
vanilla 62.12 86.62 92.52 80.42 24.35 44.78 54.36 41.16
AAT-GA COCO-CN 65.60 89.04 94.18 82.94 26.79 48.20 57.91 44.30
Flickr30k-CNA 66.00 89.40 94.44 83.28 27.59 49.16 58.79 45.18
AAT-BP COCO-CN 66.58 89.28 94.36 83.41 27.68 49.51 59.17 45.45
Flickr30k-CNA 67.18 89.92 94.56 83.89 27.99 49.81 59.49 45.76
Table 10: Recall rates on Flickr30k-CNA for text-to-image retrieval, based on ViT-B. AAT is optimized on the COCO-CN val set and a 1k-sampled subset of the Flickr30k-CNA val set, respectively.
5.3 Ablation Study
5.3.1 The Value of $\beta$
We empirically set $\beta=0.1$ in AAT-GA and study its impact on performance. Figure 3 shows the relationship between decreasing $\beta$ values and mean-R on the COCO-CN all set. As $\beta$ decreases from 1 to 0.02, mean-R initially improves, peaking at $\beta=0.1$, and then declines. For AAT-BP, we examine the $\beta$ values derived from the sigmoid function applied to the learnable $\alpha$ parameters with temperature $\tau$. We observe that $\beta$ values in deeper transformer layers tend to be more polarized—clustered near 0 or 1—whereas those in shallower layers remain closer to the initial value of 1. This suggests that gradient propagation during BP optimization varies across layers, which may help explain the performance differences between AAT-GA and AAT-BP observed in certain scenarios.
5.3.2 The Size of $\mathcal{D}$
We ablate the effect of the validation set $\mathcal{D}$ size, which originally contains 1k image-text pairs for AAT optimization. Subsets of 100, 200, and 500 image-text pairs are randomly sampled from $\mathcal{D}$, and the resulting mean-R for AAT-improved CLIP and Chinese-CLIP are reported in Table 9. Even with only 100 image-text pairs, AAT outperforms the baseline ($\text{size}=0$) by around 1% for CLIP and 3% for Chinese-CLIP. As the size increases to 500, the further performance gain becomes marginal.
5.3.3 Selection of AAT Optimization Data
As described in Section 5.1, we use a 1k-subset from MS COCO or COCO-CN to optimize AAT. To examine the impact of optimization data selection and potential domain bias, we construct a new $\mathcal{D}$ using 1k images from the Flickr30k-CNA val set, each paired with one corresponding caption. We then re-run AAT using this in-domain $\mathcal{D}$ and compare its performance on Flickr30k-CNA against that of AAT optimized with COCO-CN samples. As shown in Table 10, using in-domain data yields only marginal improvements (typically $<$0.5%), implicitly demonstrating the strong generalization of AAT-refined representations across domains.
size metric vanilla AAT-BP AAT-GA AAT-GA (optimized)
ViT-B mean-R 58.9 60.9 61.5 61.2
runtime - 2 min (0.3 GPU hours) 30 min (4 GPU hours) 18 min (2.4 GPU hours)
ViT-L mean-R 63.7 66.2 66.1 65.9
runtime - 5 min (0.7 GPU hours) 1.5 h (12 GPU hours) 27 min (3.6 GPU hours)
ViT-H mean-R 66.8 68.1 68.7 68.3
runtime - 8 min (1.1 GPU hours) 3 h (24 GPU hours) 50 min (6.7 GPU hours)
Table 11: Mean-R on the MS COCO test set for text-to-image retrieval vs. the actual AAT optimization runtime (GPU hours). “Optimized” refers to AAT-GA with optimized hyper-parameter settings.
5.4 AAT Optimization Runtime
We use 8×8$\times$8 × NVIDIA RTX 4090 GPUs and an Intel(R) Xeon(R) Platinum 8358P CPU @ 2.60GHz to optimize AAT. To demonstrate its real-world feasibility, we report the actual runtime and GPU hours under the standard configuration described in Section 5.1, as listed in the fourth and fifth columns of Table 11. While GA offers key advantages for low-precision (like INT8/INT16) inference devices, as no need for gradients during optimization, it incurs roughly 20×20$\times$20 × the time cost of BP due to the large number of fitness evaluations (even with the small validation set $\mathcal{D}$).
AAT is intentionally designed as an “out-of-the-box” practical solution, with GA hyper-parameters, such as the number of generations and crossover/mutation rates, kept consistent across different model sizes and languages to minimize manual tuning. Despite its simplicity and demonstrated effectiveness, the use of unified settings may introduce redundant computational overhead. Further hyper-parameter tuning could significantly reduce this cost without compromising performance.
Specifically, we implement the following optimizations. First, based on the ablated head analysis in Appendix C, where effective ablation ratios typically fall between 30% and 40%, we skip fitness evaluations for candidates with ablation ratios exceeding 60%. Second, we reduce the GA population size for ViT-L-based and ViT-H-based models, as smaller populations maintain comparable performance while significantly lowering runtime. Lastly, running 100 generations proves excessive: performance gains plateau during the final 30% of generations, even as fitness scores continue to improve—suggesting that over-optimization on $\mathcal{D}$ offers diminishing benefit on test data.
With these targeted adjustments, we achieve a significantly more efficient GA setup, reducing optimization overhead by 40%, 70%, and 72% for the base, large, and huge models, respectively, with only a slight drop in performance, as shown in Table 11. Detailed performance and runtime comparisons under different hyper-parameter configurations are provided in Appendix D. Notably, as a post-training technique, both AAT-BP and AAT-GA incur minimal overhead compared to training CLIP from scratch, underscoring their practicality for real-world deployment.
6 Conclusion
To further understand the role of the attention mechanism in CLIP, we conduct a comprehensive analysis of all attention heads in its image encoder. Based on the hypothesis that certain heads may negatively impact downstream performance, we propose AAT, a simple yet effective method that refines output representations by systematically ablating detrimental heads through attention weight manipulation. To identify which heads to ablate, we introduce two alternative strategies: GA and BP. AAT-GA is well-suited for scenarios with limited floating-point computation, such as inference-friendly edge devices, while AAT-BP offers greater optimization efficiency when computational resources are less constrained. Extensive experiments demonstrate that AAT is a versatile and powerful technique for enhancing CLIP-family models in diverse downstream tasks. Ultimately, we hope this study contributes to a deeper understanding of CLIP and paves the way for more efficient use of vision-language models.
Appendix A Benchmark Details and Evaluation Metrics
A.1 MS COCO
MS COCO is a large-scale dataset with 91 object categories in natural contexts, comprising about 113,287 training, 5,000 validation, and 5,000 test images (with annotations not publicly available), annotated with bounding boxes, segmentation, and 5 descriptive texts per image. For our experiments, we reconstruct the available training and validation images into two test sets: the test set, which corresponds to the original validation set, and the all set, which includes both the training and validation images. We evaluate MS COCO for text-to-image and image-to-text retrieval using recall rates at k (R@1, R@5, R@10) and mean-R (the average of R@1, R@5, and R@10) as the evaluation metrics.
A.2 Flickr30k
Flickr30k contains 31k images collected from Flickr, each accompanied by 5 reference sentences provided by human annotators. The dataset is split into 29k training images, 1k validation images, and 1k test images. Similarly to MS COCO, we treat the test images as the test set and the entire dataset as the all set. Both reconstructed sets are used to evaluate our method for text-to-image and image-to-text retrieval, with recall rates as the evaluation metric.
A.3 ReCoS
ReCoS is a benchmark designed for image-text retrieval in real-world scenarios, comprising 12 overlapping subsets with a total of 1k test images and diverse annotations per subset. We follow the standardized ReCoS-v1 evaluation protocol for experiments and report R@1 performance for text-to-image retrieval across all 12 subsets, as done in the original ReCoS paper. The subset abbreviations in Table 5 are: ACT for action, CEL for celebrity, COL for color, CNT for count, FIG for figure, OBJ for object, POS for position, SCE for scene, CR for code reasoning, NC for numerical calculation, and TT for text translation.
A.4 COCO-CN
COCO-CN is a Chinese image-text retrieval dataset with about 18k training, 1k validation, and 1k test images, annotated with 27,218 manually written Chinese sentences. In our experiments, we use the COCO-CN test images as the test set and all images as the all set, evaluating both sets with the same metric used for MS COCO.
A.5 Flickr30k-CNA
Flickr30k-CNA is a Chinese-translated version of Flickr30k, with text generated by professional English and Chinese linguists. We process and evaluate this dataset in the same manner as Flickr30k.
A.6 ImageNet-1k
ImageNet-1k is a widely used image classification benchmark originally curated for the Large Scale Visual Recognition Challenge (ILSVRC). It contains 1k object categories, with over 1.2 million labeled training images, 50,000 validation images, and 100,000 test images. ImageNet-1k also serves as a standard benchmark for zero-shot classification. During inference, object class names (provided in the CLIP repository [^1] https://github.com/OFA-Sys/Chinese-CLIP/blob/master/zeroshot_dataset_en.md) are converted into textual prompts using several predefined templates [^2] https://github.com/LAION-AI/CLIP_benchmark/blob/main/clip_benchmark/datasets/en_zeroshot_classification_templates.json. In our experiments, we evaluate zero-shot classification performance on the ImageNet-1k val set using top-1 accuracy as the metric.
A.7 Cola
Cola [29] is a compositional text-to-image retrieval benchmark designed to evaluate a model’s ability to localize and compose objects with their corresponding attributes. It includes both real-world images and synthetic 3D-rendered scenes. The task requires retrieving images that match the correct configuration of objects and attributes while rejecting distractors that contain the right components in incorrect arrangements. Cola consists of approximately 1.2k composed queries involving 168 objects and 197 attributes, distributed across roughly 30K images. It features two types of queries: single-object compositions (sourced from GQA, CLEVR, and PACO) and multi-object queries. Following the original paper, mean average precision (mAP) is used for evaluating single-object subsets, while mean accuracy is used for the multi-object subset.
Appendix B Detailed Impact of Individual Attention Heads
As described in Section 3, we investigate the impact of each individual attention head in CLIP’s image encoder on text-to-image retrieval performance. Detailed results are provided in Table 17. As shown, ablating even a single attention head can outperform the vanilla baseline, reaching up to 81.1% mean-R. Conversely, removing certain critical heads leads to a significant performance drop of up to 3.7% compared to the baseline. These results highlight the distinct roles of individual heads and suggest that carefully selecting which heads to ablate could yield further performance gains.
In the main content of the paper, we identify all negatively impactful heads—those that individually yield higher retrieval performance than the vanilla baseline when ablated—and jointly ablate them in a simple strategy referred to as naive joint head ablation (naive h-a). We extend this experiment across different model sizes for both text-to-image and image-to-text retrieval on COCO-CN. As shown in Table 12, even this straightforward method surpasses the baseline by 1.4$\sim$1.6on the test set and 1.5$\sim$2.6on the all set for text-to-image retrieval, and by 0.4$\sim$1.2on the test set and 1.8$\sim$3.7on the all set for image-to-text retrieval.
tasks text-to-image retrieval image-to-text retrieval
splits the test set the all set the test set the all set
size method R@1 R@5 R@10 mR R@1 R@5 R@10 mR R@1 R@5 R@10 mR R@1 R@5 R@10 mR
B vanilla 62.2 86.9 94.3 81.1 24.0 45.4 55.8 41.7 56.2 83.8 93.4 77.8 22.6 43.0 52.7 39.4
naive h-a 62.7 89.6 95.6 82.6 25.2 47.0 57.5 43.2 55.8 85.5 93.4 78.2 23.6 45.0 55.0 41.2
L vanilla 63.9 88.7 94.5 82.4 26.7 48.6 58.5 44.6 60.4 84.2 93.3 79.3 24.4 44.6 54.5 41.2
naive h-a 65.9 89.9 95.6 83.8 28.9 51.2 61.4 47.2 61.6 86.1 93.8 80.5 27.7 47.9 59.2 44.9
H vanilla 69.6 89.9 95.8 85.1 30.4 52.4 62.2 48.3 63.1 86.7 93.0 80.9 26.3 46.5 56.0 42.9
naive h-a 71.8 91.3 97.0 86.7 32.6 55.0 64.9 50.8 64.2 87.7 94.2 82.0 28.5 48.7 59.4 45.5
Table 12: Comparison on COCO-CN for retrieval. R@k denotes recall at rank k, and mR (mean-R) is the average of R@1, R@5, and R@10. Model sizes are denoted as B (ViT-B), L (ViT-L), and H (ViT-H). “Vanilla” refers to the baseline model, while “naive h-a” indicates the model enhanced via naive joint head ablation.
Appendix C Analysis of the Ablated Heads
C.1 Layer Distribution
To illustrate the effect of AAT, we present the distribution of ablated heads in AAT-GA across layers for ViT-B, ViT-L, and ViT-H-based models in Figure 4, covering both English and Chinese versions of CLIP. The results reveal that CLIP models tend to have greater head redundancy (i.e., more detrimental heads) in the shallow and deep layers, with fewer in the intermediate layers. In contrast, Chinese-CLIP models exhibit a more balanced ablation pattern across all layers.
01234567891011005555101010103333777755554444333333330033334444555566661111222244444444333333332222555533331111444477773333layersnumCLIPChinese-CLIP    (a) ViT-B-based AAT 0246810121416182022005555101010101515151555558888666644445555555544445555555577775555111111114444555566667777555500121212121010101088886666666699997777666655553333333388885555555555551010101066668888666666665555777766666666777766661010101066664444layersnumCLIPChinese-CLIP        (b) ViT-L-based AAT 024681012141618202224262830005555101010101515151577771010101088888888555555554444999955557777555522227777555511118888999977777777777799997777444488888888333355554444999955559999888810101010101010106666999912121212555588884444444499997777666611118888333344448888555555556666666677778888888866664444666677774444444488886666layersnumCLIPChinese-CLIP                (c) ViT-H-based AAT
Figure 4: Number of ablated heads across layers in AAT for ViT-B, ViT-L, and ViT-H. ViT-B consists of 12 layers with 12 attention heads each; ViT-L has 24 layers with 16 heads per layer; and ViT-H includes 32 layers, also with 16 heads per layer.
C.2 Statistics
We report the overall ablation ratio (calculated as the number of ablated heads divided by the total number of heads) and the average number of ablated heads per layer for different model sizes in Table 13. As shown, both CLIP and Chinese-CLIP models follow similar trends: base models have ablation ratios around 0.3, while larger models reach up to 0.4. This supports the intuition that larger models contain more redundant attention heads.
model metric ViT-B ViT-L ViT-H
CLIP ratio 0.31 0.34 0.40
avg 3.7 5.4 6.4
Chinese-CLIP ratio 0.28 0.38 0.40
avg 3.4 6.1 6.4
Table 13: Overall ablation ratio and average ablated heads per layer across different model sizes and languages. The results are averaged over three AAT-GA runs with different random seeds.
Appendix D Computational Cost Analysis
We optimize AAT-GA by eliminating redundant overhead, as described in Section 5.4 of the paper. Detailed performance and runtime comparisons under different hyper-parameter configurations are provided in Tables 14 and 15.
Table 14 reports AAT-GA performance and optimization runtime across various hyper-parameter settings. Based on these results, we identify efficient GA configurations: gen = 75 and psize = 48 for ViT-B/L models, and gen = 75 and psize = 64 for ViT-H models. As shown in Table 15, skipping fitness evaluations for high-ablation-ratio candidates further improves efficiency compared to the unified configuration. The final runtime cost and corresponding performance are summarized in Table 11 in the main paper.
CLIP-ViT-B CLIP-ViT-L CLIP-ViT-H
psize 48 32 16 96 72 48 128 96 64
gen mR T mR T mR T mR T mR T mR T mR T mR T mR T
100 61.5 30 61.1 20 60.4 10 66.1 90 66.1 66 66.0 45 68.7 180 68.8 140 68.6 90
75 61.4 22 61.1 15 60.1 8 66.0 66 65.8 50 65.9 33 68.5 140 140 100 68.4 66
50 60.9 15 60.6 10 59.9 5 65.1 45 64.9 35 64.5 24 68.1 90 67.9 66 67.6 45
Table 14: Mean-R (mR) across different CLIP model sizes for text-to-image retrieval on the MS COCO test set, along with the corresponding GA optimization runtime (T, in minutes). For clarity, “gen” refers to the number of generations, and “psize” denotes the population size.
CLIP-ViT-B CLIP-ViT-L CLIP-ViT-H
mean-R runtime mean-R runtime mean-R runtime
non-skip 61.4 22 min 65.9 33 min 68.4 1.1 h
skip 61.2 18 min 65.9 27 min 68.3 50 min
Table 15: Comparison of mean-R and runtime with and without skipping fitness evaluations. Models are evaluated on the MS COCO test set for text-to-image retrieval.
Appendix E Evaluation on Compositional Retrieval
Compositional retrieval is a challenging task that goes beyond conventional cross-modal retrieval by requiring a deeper understanding of object relationships, attributes, and reasoning—rather than relying solely on visual perception. In order to explore the potential of AAT in this setting, we evaluate AAT-improved CLIP on Cola, a recently proposed challenging benchmark designed for compositional text-to-image retrieval across both real-world and synthetic domains.
Table 16 reports the results for both vanilla and AAT-improved CLIP models. AAT improves mean accuracy on the multi-object benchmark by 2.38% with GA and 3.80% with BP. For the single-object benchmark, AAT yields mAP gains of up to 2.27% on the Cola-GQA subset and up to 1.95% on the Cola-PACO subset. However, no improvement is observed on the Cola-CLEVR subset. This limitation likely stems from the synthetic nature of CLEVR, which consists of 3D-rendered objects and requires complex reasoning capabilities that exceed CLIP’s representational capacity. As a result, CLIP—whether enhanced with AAT or not—struggles on this subset, highlighting a boundary of current vision-language models in handling highly abstract compositional reasoning.
single-object multi-object
GQA CLEVR PACO
vanilla 35.36 7.57 26.38 18.10
AAT-GA 37.63 7.54 28.33 20.48
AAT-BP 36.41 7.27 27.69 21.90
Table 16: Comparison of CLIP and AAT-improved CLIP on Cola. Results are based on ViT-B, with mAP used as evaluation metric for single-object queries and mean accuracy for multi-object queries.
Appendix F Comparison with SFT
To further validate the effectiveness and practicality of AAT, we compare it against conventional supervised finetuning (SFT). We conduct this comparison using Chinese-CLIP, where AAT demonstrates more pronounced improvements than on the English counterpart, making performance differences easier to observe. Since AAT are optimized using 1k image-text pairs from COCO-CN, we train the vanilla Chinese-CLIP on the same dataset using the contrastive supervision objective employed in CLIP, following the recommended hyper-parameters from the OpenCLIP codebase [15]. Specifically, we use a base learning rate of 5e-5 with a cosine decay schedule, the AdamW optimizer, a batch size of 256, and a weight decay of 1e-3. In experiments, we vary the number of training epochs from 1 to 64 to explore optimal performance across different model sizes. We evaluate the finetuned models on both the in-domain COCO-CN test set and the cross-domain Flickr30k-CNA test set for text-to-image retrieval. The results are presented in Figure 5.
SFT exhibits varying mean-R on COCO-CN across model sizes. It slightly outperforms AAT on ViT-B ($+$0.2%), achieves a larger gain on ViT-L ($+$2.2%), and performs comparably on ViT-H (--0.1%). However, its results are highly sensitive to the number of training epochs, likely due to overfitting or convergence to local minima under limited data. This instability is consistent across repeated trials, highlighting the challenges of SFT in data-scarce settings.
Although SFT beats AAT in some cases, AAT exhibits markedly stronger cross-domain generalization. As shown in the right column of Figure 5, AAT consistently outperforms SFT by at least 1.27% on ViT-B, 0.06% on ViT-L, and 1.37% on ViT-H. Notably, the SFT models that perform best on COCO-CN exhibit substantial performance degradation on Flickr30k-CNA, in some cases even falling below the baseline. This indicates a vulnerability of SFT when trained on limited data.
Appendix G Visualization of Samples for AAT Optimization and Its Failure Cases
For the convenience of readers, we provide a visualization of the validation set $\mathcal{D}$ used for AAT optimization in Figure 6. As discussed in Appendix E and Section 5.2.1 of the paper, AAT fails on Cola-CLEVR and three non-natural subsets from ReCoS. Representative examples from these tasks are shown in Figure 7 and Figure 8, respectively. As illustrated, there exists a significant domain gap between the test images and those in $\mathcal{D}$. Moreover, certain tasks require models to perform high-level comprehension and reasoning, which exceeds the representational capacity of CLIP.
(a) Mean-R on COCO-CN, based on ViT-B.
(b) Mean-R on Flickr30k-CNA, based on ViT-B.
(c) Mean-R on COCO-CN, based on ViT-L.
(d) Mean-R on Flickr30k-CNA, based on ViT-L.
(e) Mean-R on COCO-CN, based on ViT-H.
(f) Mean-R on Flickr30k-CNA, based on ViT-H.
Figure 5: Comparison for text-to-image retrieval among AAT-improved models, the SFT models, and the vanilla counterparts. For SFT models, mean-R across increasing training epochs are reported. Evaluation is conducted on the test sets of COCO-CN and Flickr30k-CNA for each model variant.
Figure 6: Visualization of samples from the validation set $\mathcal{D}$ used for AAT optimization, sourced from MS COCO. Each image is paired with a single text description, such as, “An old woman sits next to a large stuffed teddy bear on a bench,” which corresponds to the first image in the top row.
Figure 7: Visualization of samples from Cola-CLEVR, a subset of Cola used as a retrieval benchmark in this paper. Each image is paired with a set of descriptive attributes, such as “large,” “purple,” “rubber,” “cube,” “metal,” etc.
(a) Illustrative examples from the subset of code reasoning within ReCoS-v1. Each image is paired with a text description explaining the reasoning process behind the code presented in the image.
(b) Illustrative examples from the subset of numerical calculation within ReCoS-v1. Each image is paired with a text description that either provides answers to numerical calculations or describes the geometric shapes depicted in the image.
(c) Illustrative examples from the subset of text translation within ReCoS-v1. Each image is paired with a text description that provides the English translation of the Chinese text depicted in the image. For instance, “The English translation corresponding to the text in the text-only picture is ‘Delicious dinner’,” corresponds to the first image in the top row.
Figure 8: Visualization of non-natural images from three ReCoS-v1 subsets: code reasoning (CR), numerical calculation (NC), and text translation (TT).
$l$ $h$ R@1 R@5 R@10 mR $l$ $h$ R@1 R@5 R@10 mR $l$ $h$ R@1 R@5 R@10 mR
0 0 60.8 86.0 93.5 80.1 1 0 60.3 86.3 93.7 80.1 2 0 60.8 86.3 93.3 80.1
1 59.9 86.3 93.4 79.9 1 60.7 86.2 93.2 80.0 1 60.4 86.1 93.2 79.9
2 60.4 86.2 93.2 79.9 2 60.9 86.4 93.0 80.1 2 60.1 86.3 93.2 79.9
3 60.2 86.5 93.3 80.0 3 60.6 85.7 93.3 79.9 3 60.1 85.9 93.3 79.8
4 60.8 86.3 93.5 80.2 4 60.2 85.8 93.4 79.8 4 60.3 86.0 93.3 79.9
5 60.2 86.3 93.3 79.9 5 60.3 86.1 93.2 79.9 5 60.7 85.9 93.2 79.9
6 60.5 86.1 93.4 80.0 6 60.4 86.2 93.2 79.9 6 60.3 85.8 93.4 79.8
7 60.0 86.2 93.6 79.9 7 60.2 86.4 93.2 79.9 7 58.2 85.1 93.0 78.8
8 60.4 86.4 93.2 80.0 8 60.4 86.3 93.3 80.0 8 60.4 86.6 93.1 80.0
9 60.3 86.2 93.2 79.9 9 60.6 86.3 93.2 80.0 9 60.5 86.1 93.2 79.9
10 60.4 86.4 93.2 80.0 10 60.2 86.1 93.4 79.9 10 60.3 86.3 93.4 80.0
11 60.2 85.9 93.6 79.9 11 60.4 86.0 93.2 79.9 11 60.7 86.5 93.0 80.1
3 0 60.9 85.7 93.6 80.1 4 0 60.6 86.4 93.4 80.1 5 0 60.8 85.8 93.4 80.0
1 60.1 86.1 93.1 79.8 1 59.3 85.9 93.2 79.5 1 60.4 86.2 93.4 80.0
2 59.1 85.8 92.6 79.2 2 60.4 85.8 93.2 79.8 2 60.6 86.1 93.4 80.0
3 60.0 86.2 93.3 79.8 3 60.0 86.4 93.1 79.8 3 58.9 84.4 92.4 78.6
4 59.6 85.9 93.4 79.6 4 60.2 85.8 93.4 79.8 4 60.6 86.5 93.3 80.1
5 60.3 86.6 93.1 80.0 5 59.9 86.3 93.3 79.8 5 60.2 86.0 93.3 79.8
6 60.5 86.3 93.2 80.0 6 59.7 86.4 93.3 79.8 6 59.9 86.2 93.3 79.8
7 60.5 86.3 93.4 80.1 7 61.1 86.0 93.3 80.1 7 60.6 86.4 93.6 80.2
8 60.7 86.5 93.3 80.2 8 60.4 86.4 93.4 80.1 8 60.1 85.7 93.4 79.7
9 60.1 86.5 93.3 80.0 9 59.1 85.9 93.2 79.4 9 60.3 86.5 93.1 80.0
10 59.7 86.8 93.2 79.9 10 60.3 86.4 93.5 80.1 10 60.0 85.4 93.1 79.5
11 57.4 83.3 91.0 77.2 11 60.2 86.1 92.9 79.7 11 61.0 86.0 93.5 80.2
6 0 60.4 86.3 93.1 79.9 7 0 60.0 85.8 93.4 79.7 8 0 60.4 86.5 93.3 80.1
1 59.4 86.2 93.4 79.7 1 60.1 86.2 93.3 79.9 1 59.8 86.0 92.9 79.6
2 59.9 85.3 92.9 79.4 2 61.0 86.7 93.3 80.3 2 55.0 83.7 91.2 76.6
3 60.4 86.2 93.1 79.9 3 59.3 86.0 92.8 79.4 3 60.5 86.1 93.4 80.0
4 60.6 86.3 93.4 80.1 4 60.2 86.2 92.9 79.8 4 79.9 85.9 93.3 79.9
5 59.9 86.4 93.2 79.8 5 60.5 86.2 93.3 80.0 5 60.3 86.2 93.4 80.0
6 59.5 86.3 93.4 79.7 6 59.9 86.2 92.9 79.7 6 60.7 85.7 92.8 79.7
7 60.3 86.2 93.1 79.9 7 60.3 86.5 93.1 80.0 7 60.4 86.4 93.0 79.9
8 60.3 86.1 93.5 80.0 8 60.6 86.2 93.2 80.0 8 60.0 86.7 92.8 79.8
9 60.7 86.2 93.2 80.0 9 60.1 85.9 93.1 79.7 9 59.9 86.4 93.3 79.9
10 60.3 86.2 93.2 79.9 10 59.8 86.6 93.3 79.9 10 61.0 86.2 93.6 80.3
11 60.8 86.7 93.8 80.4 11 59.9 86.2 93.2 79.8 11 60.1 86.1 93.1 79.8
9 0 59.5 86.2 93.0 79.6 10 0 60.2 86.7 93.4 80.1 11 0 61.4 87.4 93.4 80.7
1 60.8 86.4 93.3 80.2 1 60.6 86.0 93.0 79.9 1 54.7 83.0 91.0 76.2
2 60.4 86.8 93.4 80.2 2 60.7 85.6 93.3 79.9 2 59.3 84.9 91.9 78.7
3 60.1 86.4 93.2 79.9 3 59.6 86.6 93.0 79.7 3 59.4 83.5 90.9 77.9
4 59.8 86.2 92.7 79.6 4 61.0 86.9 93.4 80.4 4 60.2 86.2 93.2 79.9
5 59.7 86.7 92.8 79.7 5 60.6 85.6 93.9 80.0 5 59.9 86.2 93.4 79.8
6 60.8 87.2 93.4 80.5 6 60.6 86.8 93.2 80.2 6 61.5 87.1 94.6 81.1
7 59.4 85.3 93.4 79.4 7 62.1 86.8 93.5 80.8 7 58.3 84.4 92.4 78.4
8 61.2 86.5 93.5 80.4 8 60.9 86.7 93.4 80.3 8 58.1 84.6 92.3 78.3
9 62.0 86.8 93.4 80.7 9 61.4 87.5 93.4 80.8 9 56.4 84.6 91.0 77.3
10 60.4 86.7 93.4 80.2 10 61.1 86.2 93.4 80.2 10 60.6 86.7 93.6 80.3
11 60.0 86.8 93.3 80.0 11 61.6 86.8 93.1 80.5 11 59.2 85.0 91.9 78.7
vanilla 60.4 86.2 93.2 79.9
[…]
Table 17: Detailed text-to-image retrieval performance on the COCO-CN val set for each individually ablated head, evaluated using the ViT-B-based Chinese-CLIP model. Here, $l$ denotes the $l$-th transformer layer of the image encoder, and $h$ represents the $h$-th attention head in the multi-head attention (MHA) module. Evaluation metrics include recall rates at k (R@k) and mean recall rate (mR). Results from the vanilla baseline are reported in the last row for reference.
Appendix H Limitations
We summarize four limitations of this work, which we leave as directions for future exploration:
Sample Selection for Optimization: It remains unclear how to optimally select the samples used for AAT optimization—specifically, which image-text pairs are most effective for refining the output representations while ensuring broad generalization.
Dependence on Optimization Data: Although AAT is designed for data-scarce scenarios, it still requires a small validation set. Head selection is inherently data-dependent, even when the dataset is indeed minimal.
Limited Validation on More Challenging Tasks: AAT has not yet been evaluated on downstream tasks beyond cross-modal alignment, such as visual question answering (VQA), object detection, or semantic segmentation. Its applicability to these more complex settings remains an open question.
Lack of Per-Instance Adaptation: AAT performs ablation globally based on the entire validation set, without per-instance customization. In practice, certain attention heads may be detrimental in some contexts but beneficial in others. This suggests a promising direction: per-image adaptive head ablation, where head selection is dynamically tailored to the input characteristics.